OpenStudy (anonymous):

Determine the line of intersection of each system of equations. help would be greatly appreciated :D a.) 2x + y + z = 7 4x +3y - 3z = 13 4x _2y +2z = 14

OpenStudy (mertsj):

Do you mean the point of intersection?

hero (hero):

No, line of intersection: 1.) 2x + y + z = 7 2.) 4x +3y - 3z = 13 3.) 4x - 2y +2z = 14 Subtract Equation 3 from Equation 2 to get the line. Then subtract Equation 1 from Equation 2 to verify. You want to eliminate x variable in both subtractions.